OpenStudy (anonymous):

In physics the mass energy equivalence is that the mass of a body is measured by its energy content. This may be written algebraically as E=mc^2, where E is energy, m is mass, and c is the speed of light in a vacuum. Transform the equation so that it will be read as a solution for mass, m. Show all the steps.Translate each real-world situations into a multi-step equation. Solve the equation, showing all steps to solve the equation. Then write an explanation of the answer to the problem.

OpenStudy (anonymous):

@tkhunny @ganeshie8

OpenStudy (anonymous):

@ikram002p @SolomonZelman

OpenStudy (ikram002p):

well u have E=mc^2 , solve for m could you do that ?:)

OpenStudy (anonymous):

no.. I know how to do all the other problems just not this one in particular

OpenStudy (ikram002p):

okk :) well if i told you that 5=4 x could u solve for x ?

OpenStudy (anonymous):

wait wrong question

OpenStudy (anonymous):

i dont understand this one

OpenStudy (anonymous):

E = mc2 Divide both sides by c^2 m = E/(c^2) idk what i did

OpenStudy (ikram002p):

yes this is correct ^_^

OpenStudy (anonymous):

yay!!!